OpenStudy (anonymous):

-4.42x+0.9=-9.070-0.432x show steps

OpenStudy (anonymous):

\[-4.42x+0.9=-9.070-0.432x \] Subract 0.9 from both sides -4.42x=-9.070-1.332 Add 9.070 to both sides 4.65x=-1.332 Divide 4.65 on both sides x=-.2864516129

OpenStudy (anonymous):

Wait I did it wrong sorry

OpenStudy (anonymous):

Subtract 0.9 from both sides -4.42x=-10.607-0.432x Add 0.432x to both sides -3.988x=-10.607 Divide -3.988 from both sides x=2.659729188

OpenStudy (anonymous):

The second one is the right one

OpenStudy (anonymous):

ok